Derive the reduction formula

integeration(x2+a2)n/2dx = x(x2+a2)n/2÷n+1+ na2/n+1×integeration (x2 +a2)n/2-1dx

Use the formula to integrate integration(x2+a2)5/2

1
Expert's answer
2021-06-03T18:26:38-0400

"\int (x^2+a^2)^{\frac52}dx=\frac{x(x^2+a^2)^{\frac52}}{6}+\frac{5a^2}{6}\int (x^2+a^2)^{\frac32}dx=\frac{x(x^2+a^2)^{\frac52}}{6}+\frac{5a^2}{6}(\frac{x(x^2+a^2)^{\frac32}}{4}+\frac{3a^2}{4}\int (x^2+a^2)^{\frac12}dx)=" "\frac{x(x^2+a^2)^{\frac52}}{6}+\frac{5a^2x(x^2+a^2)^{\frac32}}{24}+\frac{5a^4}{16}(x(x^2+a^2)^{\frac12}+a^2\ln(x+(x^2+a^2)^{\frac12}))+C."